# Overcoming Output Dimension Collapse: How Sparsity Enables Zero-shot Brain-to-Image Reconstruction at Small Data Scales

### v1

@@ Abstract
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. Achieving such reconstruction with limited training data requires generalization beyond the training set, a task known as zero-shot prediction. Despite its importance, we still lack theoretical guidelines for achieving efficient and accurate reconstruction. In this paper, we provide a theoretical analysis of two widely used models for translating brain activity to latent image features. We define the data scale as the ratio of the number of training samples to the latent feature dimensionality and characterize how each model behaves across data scales. We first show that the naive linear regression model, which uses a shared set of input variables for all outputs, suffers from “output dimension collapse” at small data scales, restricting generalization beyond the training data. We then mathematically characterize the prediction error of the sparse linear regression model by deriving formulas linking prediction error with data scale and other problem parameters. Leveraging the sparsity of the brain-to-feature mapping, this approach enables accurate zero-shot prediction even at small data scales without trapping in output dimension collapse. Our results provide a theoretical guideline for achieving zero-shot reconstruction and highlight the benefits of variable selection in brain-to-image reconstruction.
keywords:
Brain decoding , Brain-to-image reconstruction , Zero-shot prediction , Sparse linear regression
[…]
@@ 1 Introduction
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. To uncover neural representations and move toward practical applications in medicine and industry, it will be necessary to reconstruct a broad range of subjective visual experiences. Nevertheless, a fundamental constraint governs this task: the space of possible visual stimuli is vast, yet we can only collect a limited number of brain-image pairs for training. To capture the full spectrum of subjective experiences, a model must be able to predict previously unseen stimuli, namely zero-shot prediction (Shirakawa et al., 2025). Despite its importance, we still lack theoretical guidelines for achieving accurate reconstruction under these constraints.
@@
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1). A translator module converts brain activity into latent features that represent the image content, and a generator module transforms these features into reconstructed images. This paper focuses on the translator module, trained on a limited pairs of brain activity and latent features. We quantify this limitation by the data scale, the ratio of the number of training pairs to the dimensionality of the latent features. To realize zero-shot reconstruction, the translator must be able to predict valid latent features for unseen stimuli at small data scales.
@@
Figure 1: Translator–Generator pipeline. The translator converts brain activity into latent features, and the generator transforms latent features into reconstructed images. A naive multivariate linear regression model and a sparse multivariate linear regression model are commonly used as translators. Whereas a naive model uses a single shared set of input variables to predict all output dimensions, a sparse model allows each output dimension to select its own subset of input variables.
Two linear regression models are widely used as translators: naive multivariate linear regression and sparse multivariate linear regression. A naive multivariate linear regression model, exemplified by ridge regression, uses a single shared set of input variables to predict all output dimensions and has been adopted in several high-profile studies (Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). This strategy is straightforward and low-cost, yet Shirakawa et al. (2025) reports that such models fail to generalize beyond the training domain. They attribute this failure to “output dimension collapse,” in which predictions are confined to the span of the training outputs under dataset bias. In contrast, a sparse multivariate linear regression model allows each output dimension to select its own subset of input variables. Earlier work adopted sparse regression as the translator, assuming that “brain-like” latent features induce a sparse brain-to-feature mapping (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
@@
In this paper, we provide a theoretical analysis of translator prediction under small data scales. First, complementing prior work that emphasizes dataset bias (Shirakawa et al., 2025), we show that small data scales cause output dimension collapse in naive multivariate linear regression. Using the mathematically derived best prediction under this collapse, we evaluate its impact on reconstruction pipelines that employ naive translators. Second, we examine how sparsity can enable accurate prediction at small data scales by characterizing the prediction error of naive and sparse translators. Our contributions are as follows:
  • 1.
    We show that reconstruction pipelines using naive linear regression suffer from output dimension collapse at small data scales.
  • 2.
    We mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters.
  • 3.
    We demonstrate that sparse linear regression can escape output dimension collapse and achieve accurate prediction at small data scales by leveraging sparse brain-to-feature structure.
@@ 2 Preliminaries
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1). In this framework, a module called the translator converts brain activity into latent features, and a module called the generator transforms these features into reconstructed images. The translator is typically a linear regression model trained on paired brain activity and latent feature data. The generator is typically a pre-trained or pre-defined model that generates images from the predicted features. Latent features serve as an intermediate representation that links neural signals and visual content, and recent studies often utilize artificial neural network features (e.g., CNN, CLIP).
@@
In contrast, sparse multivariate linear regression allows each output value to select its subset of input variables. According to the classic taxonomy (Guyon and Elisseeff, 2003), selection strategies fall into three groups: wrapper methods (Kohavi and John, 1997), which iteratively search feature subsets by maximizing predictive performance; embedded methods (e.g., Lasso (Tibshirani, 1996), Automatic Relevance Determination (ARD) (MacKay, 1992)), which embed sparsity directly in the training objective; and filter methods (e.g., CFS (Hall, 2000), SIS (Fan and Lv, 2008) ), which rank inputs independently of the learner using simple scores (e.g., correlation, mutual information). In high-dimensional settings such as brain-to-image reconstruction, computational speed is essential, and in practice, the runtime cost scales as wrapper $\gg$ embedded $\gg$ filter.
Previous theoretical work demonstrates that sparsity is particularly beneficial when the sample size is smaller than the dimensionality of the data. For a truly $s$-sparse input-output mapping, information-theoretic bounds show that an ideal estimator can attain near-oracle prediction error once $n\gtrsim s\log(d_{\mathrm{in}}/s)$ (Wainwright, 2009a). The Lasso is the most thoroughly analyzed among practical sparse methods, and theory confirms it reaches essentially the same rate and achieves accurate prediction with $n=\mathcal{O}(s\log(d_{\mathrm{in}}))$ (Donoho, 2006; Wainwright, 2009b). The SIS filter method is also proven to recover the correct support of the mapping with high probability when $d_{\mathrm{in}}$ is exponentially large in $n$ (Fan and Lv, 2008).
3 Output dimension collapse by small data scale
Although both sparse and naive (non-sparse) models are used as the translator module, recent works increasingly rely on the naive model for simplicity and computational efficiency (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). This section focuses on the naive translator and shows a critical limitation at small data scales.
In brain-to-image reconstruction, the translator must map brain activity $\bm{x}$ onto a high-dimensional image feature space $\bm{y}$ while relying on a limited training sample. If we use a naive multivariate linear regression under such scenarios, a phenomenon “output dimension collapse” (ODC) can occur, whereby the predictions become restricted to the subspace determined by the training outputs (Shirakawa et al., 2025). This issue is especially critical for zero-shot prediction when the training subspace is low-dimensional, because the model inevitably compresses unfamiliar features into that subspace.
To understand why this collapse occurs, let us examine the predicted features of ridge regression. First, we learn a weight matrix $\hat{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ by minimizing the sum of squared errors plus a regularization term:
@@ 3 Output dimension collapse by small data scale
where $\bm{m}={X_{\mathrm{tr}}}\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}\bm{x}_{\mathrm{te}}\in\mathbb{R}^{n}$, and $m_{i}\in\mathbb{R}$ is the $i$-th element of $\bm{m}$. Clearly, the prediction $\hat{\bm{y}}_{\mathrm{te}}$ lies in the subspace spanned by the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. In other words, no matter what the test input is, the predicted output is constrained to that subspace (Figure 2A). Importantly, this property is not the familiar one that confines predictions to the span of the training inputs $\left\{\bm{x}^{(i)}\right\}_{i=1}^{n}$, but instead that it confines predictions to the span of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. This subspace is determined solely by the training outputs and does not depend on the input brain data; given the same training outputs, the same property holds regardless of measurement modality (e.g., fMRI, MEG), subject, or decoding task (e.g., perception, imagery, or illusion). Moreover, this property is not limited to ridge regression; it also holds for other naive multivariate linear regression methods, such as kernel linear regression and PLS (see A.1 for details).
@@
A key condition that leads to output dimension collapse is when the training outputs lie in a low-dimensional subspace. Shirakawa et al. (2025) showed that such a collapse occurs when the training data lack sufficient diversity and the latent features exhibit a cluster structure. In such scenarios, the predicted outputs are pulled toward the clusters in the training data, making generalization to unseen stimuli difficult.
In contrast, we focus here on a different but equally critical factor: small data scales. Even in the absence of explicit cluster structure, having $n<d_{\mathrm{out}}$ inherently forces all predictions into an at-most $n$-dimensional subspace. As $n/d_{\mathrm{out}}$ grows smaller, this low-dimensional subspace becomes more limiting, producing severe output dimension collapse. Consequently, regardless of whether the training outputs form visible clusters, small data scales alone can lead to the same fundamental difficulty in zero-shot prediction.
From another perspective, output dimension collapse due to small data scale is an inevitable consequence of reducing the number of free parameters in the model. The model must learn a weight matrix with many parameters to predict high-dimensional latent features, but the available training samples are insufficient to identify all of them. Therefore, the model must reduce the effective number of free parameters by imposing additional constraints. Naive linear regression accomplishes this by confining all predictions to the linear span of the training outputs. This implicit constraint is precisely what leads to output dimension collapse.
To formally characterize this collapse, we derive the best prediction attainable within the collapsed subspace. This best prediction establishes a lower bound for prediction error under output dimension collapse. Let $\mathcal{A}=\left\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\right\}$ denote the subspace defined by linear combinations of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. The vector $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\in\mathcal{A}$ that is closest to the true latent feature vector $\bm{y}_{\mathrm{te}}$ can be derived analytically as:
@@
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}$ (7)
where $A^{+}$ denotes the Moore-Penrose pseudo-inverse of $A$ (see A.2 for details). This expression shows that $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ is the orthogonal projection of the true latent feature vector onto the subspace $\mathcal{A}$ (Figure 2B).
Crucially, the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ provides the lower bound for prediction error in naive multivariate linear regression:
@@
The first term $\lVert\hat{\bm{y}}_{\mathrm{te}}-\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\rVert^{2}$ corresponds to the within-subspace error, and the second term $\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ corresponds to the out-of-subspace error. When $n<d_{\mathrm{out}}$, the out-of-subspace error is generally non-zero, implying an irreducible error dictated by the training outputs. This error originates from the training outputs being insufficient to cover the universe of possible images. We later verify empirically that this irreducible error is large enough to degrade the quality of reconstructed images, and actual predictions from brain activity are close to this analytic limit.

### v2

# Overcoming Output Dimension Collapse: When Sparsity Enables Zero-shot Brain-to-image Reconstruction at Small Data Scales

@@ Abstract
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. A key challenge in this task is data scarcity: a translator that maps brain activity to latent image features is trained on a limited number of brain-image pairs, making the translator a bottleneck for zero-shot reconstruction beyond the training stimuli. In this paper, we provide a theoretical analysis of two translator designs widely used in recent reconstruction pipelines: naive multivariate linear regression and sparse multivariate linear regression. We define the data scale as the ratio of the number of training samples to the latent feature dimensionality and characterize the behavior of each model across data scales. We first show that the naive linear regression model, which uses a shared set of input variables for all outputs, suffers from “output dimension collapse” at small data scales, restricting generalization beyond the training data. We then analyze sparse linear regression models in a student–teacher framework and derive expressions for the prediction error in terms of data scale and other sparsity-related parameters. Our analysis clarifies when variable selection can reduce prediction error at small data scales by exploiting the sparsity of the brain-to-feature mapping. Our findings provide quantitative guidelines for diagnosing output dimension collapse and for designing effective translators and feature representations for zero-shot reconstruction.
@@ 1 Introduction
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. To uncover neural representations and move toward practical applications in medicine and industry, we need reconstruction methods that generalize to a broad range of subjective visual experiences. A key challenge in this task is data scarcity: the space of possible visual stimuli is vast, yet we can only collect a limited number of brain-image pairs for training. To capture the full spectrum of subjective experiences, a model must be able to predict previously unseen stimuli, i.e., zero-shot prediction (Shirakawa et al., 2025). Despite its importance, we still lack theoretical guidelines for achieving accurate reconstruction under these constraints.
@@
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Fig. 1). In this framework, the translator converts brain activity into latent features that represent the visual content, and the generator transforms these features into reconstructed images. Latent features are typically high-dimensional to represent a broad spectrum of visual stimuli. Recent pipelines often leverage powerful generators, but the translator must be trained using only a limited number of brain-feature pairs. We quantify this data limitation using the data scale $n/d_{\mathrm{out}}$, where $n$ is the number of training pairs and $d_{\mathrm{out}}$ is the latent feature dimensionality. The translator becomes a key bottleneck for zero-shot reconstruction because it must predict valid latent features for unseen stimuli when $n/d_{\mathrm{out}}<1$.
A common translator design is what we call naive multivariate linear regression, which uses a shared set of input variables to predict all output dimensions (e.g., ridge regression). This design is straightforward and low-cost, making it a common choice in recent studies (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). Another approach is sparse multivariate linear regression, which performs variable selection for each output dimension (e.g., Lasso, ARD, filter methods). Earlier work adopted sparse regression as the translator, assuming that “brain-like” latent features induce a sparse brain-to-feature mapping (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019). Theoretical results suggest that sparse regression can be sample-efficient when the underlying mapping is sparse (Donoho, 2006; Wainwright, 2009b). Notably, these linear translators are not only classical baselines but are still central components in many recent reconstruction pipelines.
@@
Figure 1: Translator–Generator pipeline. The translator converts brain activity into latent features, and the generator transforms latent features into reconstructed images. A naive multivariate linear regression model and a sparse multivariate linear regression model are commonly used as translators. A naive model uses a shared set of input variables to predict all output dimensions, and a sparse model performs variable selection for each output dimension.
In brain-to-image reconstruction, Shirakawa et al. (2025) pointed out a phenomenon in naive multivariate linear translators called “output dimension collapse” (ODC). They reported that several high-profile reconstruction methods fail to generalize beyond the training domain. They attributed this failure to ODC, in which predictions are confined to the span of the training outputs in latent feature space under dataset bias. ODC is a challenge that must be addressed to achieve zero-shot prediction, but we still lack insight into when it occurs and how it can be overcome.
@@
In this paper, we provide a theoretical analysis of translator prediction under small data scales. First, complementing prior work that emphasizes dataset bias (Shirakawa et al., 2025), we show that small data scales can cause ODC in naive multivariate linear regression. Using the mathematically derived best prediction under this collapse, we evaluate its impact on reconstruction pipelines that employ naive translators. Second, we examine when sparsity can enable accurate prediction at small data scales by mathematically characterizing the prediction error of naive and sparse translators. The main contributions of this work are threefold: (1) a theoretical demonstration that naive multivariate linear regression exhibits output dimension collapse at small data scales; (2) an analytic diagnostic of the practical impact of ODC via the best prediction under collapse; and (3) a characterization of when sparsity enables accurate prediction at small data scales through quantitative error predictions for naive and sparse translators.
@@ 2 Preliminaries
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Fig. 1). In this framework, a module called the translator converts brain activity into latent features, and the generator transforms these features into reconstructed images. The translator is typically a linear regression model trained on paired brain activity and latent feature data. The generator is typically a pre-trained or pre-defined model that generates images from the predicted features. Latent features serve as an intermediate representation that links neural signals and visual content, and recent studies often utilize artificial neural network features, such as Convolutional Neural Networks (CNNs) and Contrastive Language-Image Pretraining (CLIP).
@@
Another approach is sparse multivariate linear regression, which performs variable selection for each output dimension. According to the classic taxonomy (Guyon and Elisseeff, 2003), selection strategies fall into three groups: wrapper methods (Kohavi and John, 1997), which iteratively search feature subsets by maximizing predictive performance; embedded methods (e.g., Lasso (Tibshirani, 1996), Automatic Relevance Determination (ARD) (MacKay, 1992)), which embed sparsity directly in the training objective; and filter methods (e.g., CFS (Hall, 2000), SIS (Fan and Lv, 2008) ), which rank inputs independently of the learner using simple scores (e.g., correlation, mutual information). In high-dimensional settings such as brain-to-image reconstruction, computational speed is essential, and wrapper methods are typically the most computationally expensive, embedded methods have intermediate cost, and filter methods are usually the least expensive.
Previous theoretical work demonstrates that sparsity is particularly beneficial when the sample size is smaller than the dimensionality of the data. We call an input–output mapping $s$-sparse if, for each output coordinate, the corresponding weight vector has at most $s$ non-zero entries (i.e., the output depends on only $s$ of the $d_{\mathrm{in}}$ input variables). For a truly $s$-sparse input-output mapping, information-theoretic bounds show that an ideal estimator can attain near-oracle prediction error once $n\gtrsim s\log(d_{\mathrm{in}}/s)$ (Wainwright, 2009a). The Lasso is the most thoroughly analyzed among practical sparse methods, and theory confirms it reaches essentially the same rate and achieves accurate prediction with $n=\mathcal{O}(s\log(d_{\mathrm{in}}))$ (Donoho, 2006; Wainwright, 2009b). The SIS filter method is also proven to recover the correct support of the mapping with high probability when $d_{\mathrm{in}}$ grows exponentially in $n$ (Fan and Lv, 2008).
3 Output dimension collapse by small data scales
Although both naive and sparse models are used as the translator module, recent works increasingly rely on the naive model for simplicity and computational efficiency (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). This section focuses on the naive translator and shows a critical limitation at small data scales.
Output dimension collapse (ODC) refers to a phenomenon in which predictions of naive multivariate linear regression are confined to a low-dimensional subspace spanned by the training outputs in the latent feature space. This restriction is especially critical to zero-shot prediction when the training-output subspace is low-dimensional, because unfamiliar features are inevitably compressed into that subspace. To understand why this collapse occurs, let us examine the predicted features of ridge regression. In ridge regression, the fitted weight matrix $\hat{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ admits the following closed-form solution:
@@ 3 Output dimension collapse by small data scales
where $\bm{m}={X_{\mathrm{tr}}}\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}\bm{x}_{\mathrm{te}}\in\mathbb{R}^{n}$, and $m_{i}\in\mathbb{R}$ is the $i$-th element of $\bm{m}$. Clearly, the prediction $\hat{\bm{y}}_{\mathrm{te}}$ lies in the subspace spanned by the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. In other words, no matter what the test input is, the predicted output is constrained to that subspace (Fig. 2A). Importantly, this property is not the familiar one that confines predictions to the span of the training inputs $\left\{\bm{x}^{(i)}\right\}_{i=1}^{n}$, but instead that it confines predictions to the span of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. This subspace is determined solely by the training outputs and does not depend on the input brain data; given the same training outputs, the same property holds regardless of measurement modality (e.g., fMRI, MEG), subject, or decoding task (e.g., perception, imagery, or illusion). Moreover, this property is not limited to ridge regression; it also holds for other naive multivariate linear regression methods, such as kernel linear regression and PLS (see Appendix B.1 for details).
@@
A key condition that leads to ODC is when the training outputs lie in a low-dimensional subspace. Shirakawa et al. (2025) showed that such a collapse occurs when the training data lack sufficient diversity and the latent features exhibit a cluster structure. In such scenarios, the predicted outputs are pulled toward the clusters in the training data, making generalization to unseen stimuli difficult.
In contrast, we focus here on a different but equally critical factor: small data scales. Even in the absence of explicit cluster structure, having $n<d_{\mathrm{out}}$ inherently forces all predictions into an at-most $n$-dimensional subspace. As $n/d_{\mathrm{out}}$ grows smaller, this low-dimensional subspace becomes more limiting, producing severe ODC. Consequently, regardless of whether the training outputs form visible clusters, small data scales alone can lead to the same fundamental difficulty in zero-shot prediction.
From another perspective, ODC by small data scales is an inevitable consequence of reducing the number of free parameters in the model. The model must learn a weight matrix with many parameters to predict high-dimensional latent features, but the available training samples are insufficient to identify all of them. Therefore, the model must reduce the effective number of free parameters by imposing additional constraints. Naive linear regression accomplishes this by confining all predictions to the linear span of the training outputs. This implicit constraint is precisely what leads to ODC.
To formally characterize this collapse, we derive the best prediction attainable within the collapsed subspace. This best prediction establishes a lower bound for prediction error under ODC. Let $\mathcal{A}=\left\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\right\}$ denote the subspace defined by linear combinations of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. The vector $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\in\mathcal{A}$ that is closest to the true latent feature vector $\bm{y}_{\mathrm{te}}$ can be derived analytically as:
@@
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}$ (7)
where $A^{\dagger}$ denotes the Moore-Penrose pseudo-inverse of $A$ (see Appendix B.2 for details). This expression shows that $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ is the orthogonal projection of the true latent feature vector onto the subspace $\mathcal{A}$ (Fig. 2B). Because it represents the best achievable prediction within this subspace, computing it in a given reconstruction pipeline allows us to assess how strongly ODC constrains its predictions directly.
The best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ provides the lower bound for prediction error in naive multivariate linear regression:
@@
The first term $\lVert\hat{\bm{y}}_{\mathrm{te}}-\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\rVert^{2}$ corresponds to the within-subspace error, and the second term $\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ corresponds to the out-of-subspace error. When $n<d_{\mathrm{out}}$, the out-of-subspace error is generally non-zero, implying an irreducible error dictated by the training outputs. This error originates from the training outputs being insufficient to cover the universe of possible images. We later verify empirically that this irreducible error is large enough to degrade the quality of reconstructed images, and most of the observed prediction error can be explained by this irreducible out-of-subspace component.